# Adversarial Robustness in Two-Stage Learning-to-Defer: Algorithms and Guarantees

### v1

@@ Abstract
Learning-to-Defer (L2D) facilitates optimal task allocation between AI systems and decision-makers. Despite its potential, we show that current two-stage L2D frameworks are highly vulnerable to adversarial attacks, which can misdirect queries or overwhelm decision agents, significantly degrading system performance. This paper conducts the first comprehensive analysis of adversarial robustness in two-stage L2D frameworks. We introduce two novel attack strategies—untargeted and targeted—that exploit inherent structural vulnerabilities in these systems. To mitigate these threats, we propose SARD, a robust, convex, deferral algorithm rooted in Bayes and $(\mathcal{R},\mathcal{G})$-consistency. Our approach guarantees optimal task allocation under adversarial perturbations for all surrogates in the cross-entropy family. Extensive experiments on classification, regression, and multi-task benchmarks validate the robustness of SARD.
@@ 1 Introduction
Learning-to-Defer (L2D) is a powerful framework that enables decision-making systems to optimally allocate queries among multiple agents, such as AI models, human experts, or other decision-makers (Madras et al., 2018). In the two-stage framework, agents are trained offline to harness their domain-specific expertise, enabling the allocation of each task to the decision-maker with the highest confidence (Mao et al., 2023a, 2024d; Montreuil et al., 2024). L2D is particularly valuable in high-stakes applications where reliability and performance are critical (Mozannar and Sontag, 2020; Verma et al., 2022). In healthcare, L2D systems integrate an AI diagnostic model with human specialists, delegating routine tasks to AI while deferring edge cases, such as anomalous imaging data, to specialists for nuanced evaluation (Mozannar et al., 2023). By dynamically assigning tasks to the most suitable agent, L2D ensures both accuracy and reliability, making it ideal for safety-critical domains.
Robustness in handling critical decisions is, therefore, essential for such systems. However, existing L2D frameworks are typically designed under the assumption of clean, non-adversarial input data, leaving them highly susceptible to adversarial perturbations—subtle input manipulations that disrupt task allocation and alter decision boundaries. Unlike traditional machine learning systems, where adversarial attacks primarily affect prediction outputs (Goodfellow et al., 2014; Szegedy et al., 2014; Awasthi et al., 2023), L2D systems are susceptible to more sophisticated adversarial threats, including query redirection to less reliable agents or intentional agent overloading. These vulnerabilities severely impact performance, drive up operational costs, and compromise trust.
This paper addresses the critical yet unexplored challenge of adversarial robustness in L2D systems. Inspired by adversarial attacks on classification (Goodfellow et al., 2014; Madry et al., 2017; Gowal et al., 2020), we introduce two novel attack strategies tailored to two-stage L2D: untargeted attacks, which disrupt agent allocation, and targeted attacks, which redirect queries to specific agents. To counter these attacks, we propose a robust family of surrogate losses based on cross-entropy (Mao et al., 2023a, 2024d; Montreuil et al., 2024), designed to ensure robustness in classification, regression, and multi-task settings. Building upon advances in consistency theory for adversarial robustness (Bao et al., 2021; Awasthi et al., 2022, 2023; Mao et al., 2023b), we establish both Bayes-consistency and $(\mathcal{R},\mathcal{G})$ consistency for our surrogate losses, enabling reliable task allocation even under adversarial scenarios. Our algorithm, SARD, leverages these guarantees while preserving convexity.
@@
    We introduce two novel adversarial attack strategies for two-stage L2D: targeted attacks that redirect queries and untargeted attacks that disrupt task allocation.
@@
    We propose a robust family of surrogate losses with guarantees of Bayes-consistency and $(\mathcal{R},\mathcal{G})$ consistency across classification, regression, and multi-task settings. Our convex algorithm, SARD, leverages these guarantees to achieve robust task allocation.
@@
    We empirically demonstrate that our novel attacks effectively exploit vulnerabilities in state-of-the-art two-stage L2D approaches, while our algorithm, SARD, exhibits strong robustness against the attack across diverse tasks.
This work lays the first theoretical foundation for adversarial robustness in L2D systems.
@@ 2 Related Works
Learning-to-Defer:
The first single-stage L2D approach was introduced by Madras et al. (2018), training both the predictor and the rejector, which were built upon the framework established in Cortes et al. (2016). In a seminal work, Mozannar and Sontag (2020) proposed the first approach proven to be Bayes-consistent, ensuring optimal allocation. Verma et al. (2022) presented an alternative formulation based on one-versus-all surrogates, also proven to be Bayes-consistent, and later extended to a broader family of losses by Charusaie et al. (2022). More recently, Cao et al. (2024) proposed an asymmetric softmax surrogate to improve probability estimation between agents, addressing limitations in both Mozannar and Sontag (2020) and Verma et al. (2022). Furthermore, Mozannar et al. (2023) demonstrated that the approaches from Mozannar and Sontag (2020); Verma et al. (2022) are not $-\mathcal{H}$-consistent, leading to suboptimal performance for some distributions. Mao et al. (2024b) generalized the work of Mozannar and Sontag (2020) proving both Bayes and $\mathcal{H}$-consistency, while Mao et al. (2024c) extended it to $-\mathcal{H}$-consistency.
In the two-stage setting, where agents are already trained offline, Mao et al. (2023a) introduced the first classification approach that guarantees both Bayes-consistency and $\mathcal{H}$-consistency. This work was further extended by Mao et al. (2024d), who adapted the two-stage framework to regression tasks while maintaining these consistency guarantees. Additionally, Montreuil et al. (2024) generalized the approach to multi-task learning.
@@ 3 Preliminaries
We consider a multi-task setting that addresses both classification and regression problems simultaneously. Let $\mathcal{X}$ denote the input space, $\mathcal{Y}=\{1,\ldots,n\}$ represent the set of $n$ distinct classes for classification, and $\mathcal{T}\subseteq\mathbb{R}$ denote the target space for regression. Each data point is represented as a triplet $z=(x,y,t)\in\mathcal{Z}$, where $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}\times\mathcal{T}$. We assume the data is drawn independently and identically distributed (i.i.d.) from an underlying distribution $\mathcal{D}$ over $\mathcal{Z}$. To model this multi-task problem, we introduce a backbone $w\in\mathcal{W}$, which acts as a shared feature extractor. The backbone maps inputs $x\in\mathcal{X}$ to a latent feature representation $q\in\mathcal{Q}$, via the function $w:\mathcal{X}\to\mathcal{Q}$. Building upon this backbone, we define a classifier $h\in\mathcal{H}$, representing all possible classification heads. Formally, $h:\mathcal{Q}\times\mathcal{Y}\to\mathbb{R}$ is a scoring function, with predictions computed as $h(q)=\arg\max_{y\in\mathcal{Y}}h(q,y)$. Similarly, we define a regressor $f\in\mathcal{F}$, which maps latent features to real-valued targets, $f:\mathcal{Q}\to\mathcal{T}$. These components are integrated into a single multi-head network $g\in\mathcal{G}$, defined as $\mathcal{G}=\{g:g(x)=(h\circ w(x),f\circ w(x))\mid w\in\mathcal{W},h\in\mathcal{H},f\in\mathcal{F}\}$.
Consistency in classification:
In classification, the primary objective is to identify a classifier $h\in\mathcal{H}$ that minimizes the true error $\mathcal{E}_{\ell_{01}}(h)$, defined as $\mathcal{E}_{\ell_{01}}(h)=\mathbb{E}_{(x,y)}[\ell_{01}(h,x,y)]$. The Bayes-optimal error is expressed as $\mathcal{E}^{B}_{\ell_{01}}(\mathcal{H})=\inf_{h\in\mathcal{H}}\mathcal{E}_{\ell_{01}}(h)$. However, minimizing $\mathcal{E}_{\ell_{01}}(h)$ directly is challenging due to the non-differentiability of the true multiclass 0-1 loss (Zhang, 2002; Steinwart, 2007; Awasthi et al., 2022). To address this challenge, surrogate losses are employed as convex, non-negative upper bounds on $\ell_{01}$. A notable family of multiclass surrogate losses is the comp-sum (Mohri et al., 2012; Mao et al., 2023b), which we refer to as a family of multiclass surrogate losses:
@@
$\Psi^{u}(v)=\begin{cases}\log(1+v)&\text{if }u=1,\\ \frac{1}{1-u}\left[(1-v)^{1-u}-1\right]&\text{if }u>0\land u\neq 1.\end{cases}$ (2)
This formulation generalizes several well-known loss functions, including the sum-exponential loss (Weston and Watkins, 1998), logistic loss (Ohn Aldrich, 1997), generalized cross-entropy (Zhang and Sabuncu, 2018), and mean absolute error loss (Ghosh et al., 2017). The corresponding true error for $\Phi_{01}^{u}$ is defined as $\mathcal{E}_{\Phi_{01}^{u}}(h)=\mathbb{E}_{(x,y)}[\Phi_{01}^{u}(h,x,y)]$, with its optimal value expressed as $\mathcal{E}^{\ast}_{\Phi_{01}^{u}}(\mathcal{H})=\inf_{h\in\mathcal{H}}\mathcal{E}_{\Phi_{01}^{u}}(h)$.
A key property of a surrogate loss is Bayes-consistency, which ensures that minimizing the surrogate excess risk leads to minimizing the true excess risk (Zhang, 2002; Bartlett et al., 2006; Steinwart, 2007; Tewari and Bartlett, 2007). Formally, $\Phi_{01}^{u}$ is Bayes-consistent with respect to $\ell_{01}$ if, for any sequence $\{h_{k}\}_{k\in\mathbb{N}}\subset\mathcal{H}$, the following implication holds:
@@
This property typically assumes $\mathcal{H}=\mathcal{H}_{\text{all}}$, which may not hold for restricted hypothesis classes such as $\mathcal{H}_{\text{lin}}$ or $\mathcal{H}_{\text{ReLU}}$ (Long and Servedio, 2013; Awasthi et al., 2022; Mao et al., 2024a). To characterize consistency with a particular hypothesis set, Awasthi et al. (2022) introduced $\mathcal{H}$-consistency bounds, which rely on a non-decreasing function $\Gamma:\mathbb{R}^{+}\to\mathbb{R}^{+}$ and take the following form:
@@
Adversarial robustness:
Adversarial robust classification aims to train classifiers that are robust to small, imperceptible perturbations of the input (Goodfellow et al., 2014; Madry et al., 2017). The objective is to minimize the true multiclass loss $\ell_{01}$ evaluated on an adversarial input $x^{\prime}=x+\delta$ (Gowal et al., 2020; Awasthi et al., 2022). A perturbation $\delta$ is constrained by its magnitude, and we define the adversarial region around $x$ as $B_{p}(x,\gamma)=\{x^{\prime}\mid\|x^{\prime}-x\|_{p}\leq\gamma\}$, where $\|\cdot\|_{p}$ is the $p$-norm and $\gamma\in(0,1)$ specifies the maximum allowed perturbation. The adversarial true multiclass loss ~$\widetilde{\ell}_{01}:\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\to\{0,1\}$ is given by:
~$\widetilde{\ell}_{01}(h,x,y)=\sup_{x^{\prime}\in B_{p}(x,\gamma)}\ell_{01}(h(x^{\prime}),y).$ (5)
Similarly to classification, minimizing ~$\widetilde{\ell}_{01}$ is computationally infeasible (Zhang and Agarwal, 2020; Bartlett et al., 2006; Awasthi et al., 2022). To address this, we introduce the family of adversarial margin surrogate losses $\widetilde{\Phi}^{\rho,u}_{01}$ from the comp-sum $\rho$-margin family, which approximate the adversarial true multiclass loss ~$\widetilde{\ell}_{01}$. This family is defined as:
@@
Here, $\Psi^{u}$ and $\Psi_{\rho}$ represent transformations that characterize the behavior of the family, where the non-convex transformation is defined as $\Psi_{\rho}(v)=\min\left\{\max\left(0,1-\frac{v}{\rho}\right),1\right\}$. Recent studies have demonstrated that algorithms employing smooth regularized variants of the comp-sum $\rho$-margin losses achieve $\mathcal{H}$-consistency, thereby offering strong theoretical guarantees (Awasthi et al., 2022, 2023; Mao et al., 2023b).
Two-stage Learning-to-Defer:
The Learning-to-Defer framework assigns queries $x\in\mathcal{X}$ to the most confident agent, aiming to enhance performance by leveraging the strengths of multiple agents. The agents consist of a primary model and $J$ experts, denoted by the set $\mathcal{A}=\{0\}\cup[J]$, where 00 corresponds to the primary model $g$ defined in Section 3. Each expert $\text{M}_{j}$ provides a prediction pair $m_{j}(x)=(m_{j}^{h}(x),m_{j}^{f}(x))$, where $m_{j}^{h}(x)\in\mathcal{Y}$ is a categorical prediction and $m_{j}^{f}(x)\in\mathcal{T}$ is a regression estimate. The combined predictions of all $J$ experts are represented as $m(x)=\big{(}m_{1}(x),\dots,m_{J}(x)\big{)}$, which lies in the joint prediction space $\mathcal{M}$. In the two-stage setting, all agents are trained offline, and the framework focuses on query allocation, keeping agent parameters fixed.
A rejector function $r\in\mathcal{R}$, defined as $r:\mathcal{X}\times\mathcal{A}\to\mathbb{R}$, is learned to assign a query $x$ to the agent $j\in\mathcal{A}$ with the highest rejection score $r(x)=\arg\max_{j\in\mathcal{A}}r(x,j)$, as described by Mao et al. (2024d, 2023a); Montreuil et al. (2024).
Definition 1 (Two-Stage L2D losses).
Let an input $x\in\mathcal{X}$, for any $r\in\mathcal{R}$, we have the true deferral loss:
$\ell_{\text{def}}(r,g,m,z)$ $=\sum_{j=0}^{J}c_{j}(g(x),m_{j}(x),z)1_{r(x)=j},$
and its family of convex, non-negative, upper-bound surrogate deferral losses:
@@
where $c_{j}$ denotes the non-negative bounded cost of assigning a decision to agent $j\in\mathcal{A}$ (Madras et al., 2018). If the rejector $r\in\mathcal{R}$ assigns $r(x)=0$, the query is handled by the primary model $g$, which predicts $g(x)=(h(w(x)),f(w(x)))$ and incurs a general cost $c_{0}(g(x),z)=\psi(g(x),z)$. Here, $\psi:\mathcal{Y}\times\mathcal{T}\times\mathcal{Z}\to\mathbb{R}^{+}$ is a general measure used to quantify the prediction quality of $g$ with respect to $z$. If $r(x)=j$ for some $j>0$, the query is deferred to expert $j$, incurring a cost $c_{j}(m_{j}(x),z)=\psi(m_{j}(x),z)+\beta_{j}$, where $\beta_{j}$ represents the consultation cost associated with expert $j$. The aggregated cost across all agents is then defined as:
$\tau_{j}(g(x),m(x),z)=\sum_{i=0}^{J}c_{i}(g(x),m_{i}(x),z)1_{i\neq j}$ (7)
@@
recovering the formulation from Mao et al. (2024d) and Montreuil et al. (2024). Note that in the case of classification, the function $\psi$ corresponds to the $\ell_{01}$ loss.
4 Adversarial Attacks on Two-Stage L2D
Motivation and Setting

### v2

@@ Abstract
Two-stage Learning-to-Defer (L2D) enables optimal task delegation by assigning each input to either a fixed main model or one of several offline experts, supporting reliable decision-making in complex, multi-agent environments. However, existing L2D frameworks assume clean inputs and are vulnerable to adversarial perturbations that can manipulate query allocation—causing costly misrouting or expert overload. We present the first comprehensive study of adversarial robustness in two-stage L2D systems. We introduce two novel attack strategies—untargeted and targeted—which respectively disrupt optimal allocations or force queries to specific agents. To defend against such threats, we propose SARD, a convex learning algorithm built on a family of surrogate losses that are provably Bayes-consistent and $(\mathcal{R},\mathcal{G})$-consistent. These guarantees hold across classification, regression, and multi-task settings. Empirical results demonstrate that SARD significantly improves robustness under adversarial attacks while maintaining strong clean performance, marking a critical step toward secure and trustworthy L2D deployment.
Learning-to-Defer, Learning to reject, routing, robustness, Statistical Learning, Machine Learning
@@ 1 Introduction
Learning-to-Defer (L2D) is a powerful framework that enables decision-making systems to optimally allocate queries among multiple agents, such as AI models, human experts, or other decision-makers (Madras et al., 2018). In the two-stage framework, agents are trained offline to harness their domain-specific expertise, enabling the allocation of each task to the decision-maker with the highest confidence (Mao et al., 2023a, 2024c; Montreuil et al., 2024). L2D is particularly valuable in high-stakes applications where reliability and performance are critical (Mozannar & Sontag, 2020; Verma et al., 2022). In healthcare, L2D systems integrate an AI diagnostic model with human specialists, delegating routine tasks to AI and deferring edge cases—such as difficult images—to specialists for nuanced evaluation (Mozannar et al., 2023; Strong et al., 2024). By dynamically assigning tasks to the most suitable agent, L2D ensures both accuracy and reliability, making it ideal for safety-critical domains.
Robustness in handling high-stakes decisions is, therefore, essential for such systems. However, existing L2D frameworks are typically designed under the assumption of clean, non-adversarial input data, leaving them highly susceptible to adversarial perturbations—subtle input manipulations that disrupt task allocation and alter decision boundaries. Unlike traditional machine learning systems, where adversarial attacks primarily affect prediction outputs (Goodfellow et al., 2014; Szegedy et al., 2014; Awasthi et al., 2023), L2D systems are susceptible to more sophisticated adversarial threats, including query redirection to less reliable agents or intentional agent overloading. These vulnerabilities severely impact performance, drive up operational costs, and compromise trust.
This paper addresses the important yet unexplored challenge of adversarial robustness in L2D systems. Inspired by adversarial attacks on classification (Goodfellow et al., 2014; Madry et al., 2017; Gowal et al., 2020), we introduce two novel attack strategies tailored to two-stage L2D: untargeted attacks, which disrupt agent allocation, and targeted attacks, which redirect queries to specific agents. To counter these attacks, we propose a robust family of surrogate losses based on cross-entropy (Mao et al., 2023a, 2024c; Montreuil et al., 2024), designed to ensure robustness in classification, regression, and multi-task settings. Building upon advances in consistency theory for adversarial robustness (Bao et al., 2020; Awasthi et al., 2022, 2023; Mao et al., 2023b), we establish both Bayes-consistency and $(\mathcal{R},\mathcal{G})$ consistency for our surrogate losses, enabling reliable task allocation even under adversarial scenarios. Our algorithm, SARD, leverages these guarantees while preserving convexity.
@@
    We introduce two new adversarial attack strategies tailored to two-stage L2D: targeted attacks that manipulate routing toward specific agents, and untargeted attacks that disrupt optimal allocation.
@@
    We propose a robust family of cross-entropy-based surrogate losses that are Bayes-consistent and $(\mathcal{R},\mathcal{G})$-consistent, enabling reliable allocation in classification, regression, and multi-task settings. These losses are the foundation of our convex and efficient algorithm, SARD.
@@
    We empirically demonstrate that our attacks reveal critical vulnerabilities in state-of-the-art two-stage L2D systems, while our proposed method SARD consistently withstands these attacks and achieves robust performance across diverse tasks.
This work lays the theoretical and algorithmic foundations for adversarially robust Learning-to-Defer systems.
@@ 2 Related Works
Learning-to-Defer.
Learning-to-Defer addresses the problem of allocating decisions between an automated model and offline decision-makers. In contrast to classical selective prediction—which simply rejects uncertain inputs without further action (Chow, 1970; Bartlett & Wegkamp, 2008; Cortes et al., 2016; Geifman & El-Yaniv, 2017)—L2D explicitly learns to defer to experts based on their distribution. Existing approaches fall into two broad categories: one-stage methods, where the predictor and deferral mechanism are trained jointly, and two-stage methods, where the deferral policy is optimized post hoc, assuming access to frozen predictions from pre-trained agents. This paper focuses on the two-stage setting.
The one-stage formulation was initiated by Madras et al. (2018), who jointly trained a predictor and a rejector in the binary setting. A major theoretical milestone was achieved by Mozannar & Sontag (2020), who established the first Bayes-consistent L2D method for multiclass classification, ensuring asymptotically optimal delegation. This was followed by Verma et al. (2022), who introduced a one-versus-all surrogate formulation that also achieves Bayes-consistency and was later generalized to broader surrogate families by Charusaie et al. (2022). To improve calibration across agents, Cao et al. (2024) proposed an asymmetric softmax surrogate that addresses the shortcomings identified in both Mozannar & Sontag (2020) and Verma et al. (2022). However, Mozannar et al. (2023) showed that these methods are not $-\mathcal{H}$-consistent, potentially degrading performance when the hypothesis class is restricted. This issue was resolved by Mao et al. (2024b), who extended the theory to cover the realizable setting.
In the two-stage regime, where experts and the main model are fixed and trained offline, Narasimhan et al. (2022) introduced a post-hoc deferral strategy. Building on this, Mao et al. (2023a) provided the first Bayes- and $\mathcal{H}$-consistent formulation for classification, which was later extended to regression by Mao et al. (2024c). Finally, Montreuil et al. (2024) generalized the two-stage framework to multi-task learning scenarios.
@@ 3 Preliminaries
We consider a multi-task setting that addresses both classification and regression problems simultaneously. Let $\mathcal{X}$ denote the input space, $\mathcal{Y}=\{1,\ldots,n\}$ represent the set of $n$ distinct classes for classification, and $\mathcal{T}\subseteq\mathbb{R}$ denote the target space for regression. Each data point is represented by a triplet $z=(x,y,t)\in\mathcal{Z}$, where $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}\times\mathcal{T}$. We assume the data is drawn independently and identically distributed (i.i.d.) from an underlying distribution $\mathcal{D}$ over $\mathcal{Z}$. To model this multi-task problem, we introduce a backbone $w\in\mathcal{W}$, which acts as a shared feature extractor. The backbone maps inputs $x\in\mathcal{X}$ to a latent feature representation $q\in\mathcal{Q}$, via the function $w:\mathcal{X}\to\mathcal{Q}$. Building upon this backbone, we define a classifier $h\in\mathcal{H}$, representing all possible classification heads. Formally, $h:\mathcal{Q}\times\mathcal{Y}\to\mathbb{R}$ is a scoring function, with predictions computed as $h(q)=\arg\max_{y\in\mathcal{Y}}h(q,y)$. Similarly, we define a regressor $f\in\mathcal{F}$, which maps latent features to real-valued targets, $f:\mathcal{Q}\to\mathcal{T}$. These components are integrated into a single multi-head network $g\in\mathcal{G}$, defined as $\mathcal{G}=\{g:g(x)=(h\circ w(x),f\circ w(x))\mid w\in\mathcal{W},h\in\mathcal{H},f\in\mathcal{F}\}$.
Consistency in classification.
In classification, the primary objective is to identify a classifier $h\in\mathcal{H}$ in the specific case where $w(x)$ is the identity function, such that $h(x)=\operatorname*{arg\,max}_{y\in\mathcal{Y}}h(x,y)$. This classifier should minimize the true error $\mathcal{E}_{\ell_{01}}(h)$, defined as $\mathcal{E}_{\ell_{01}}(h)=\mathbb{E}_{(x,y)}[\ell_{01}(h,x,y)]$. The Bayes-optimal error is expressed as $\mathcal{E}^{B}_{\ell_{01}}(\mathcal{H})=\inf_{h\in\mathcal{H}}\mathcal{E}_{\ell_{01}}(h)$. However, minimizing $\mathcal{E}_{\ell_{01}}(h)$ directly is challenging due to the non-differentiability of the true multiclass 0-1 loss (Zhang, 2002; Steinwart, 2007; Awasthi et al., 2022). To address this challenge, surrogate losses are employed as convex, non-negative upper bounds on $\ell_{01}$. A notable family of multiclass surrogate losses is the comp-sum (Mohri et al., 2012; Mao et al., 2023b), which we refer to as a family of multiclass surrogate losses:
@@
[…]
This formulation generalizes several well-known loss functions, including the sum-exponential loss (Weston & Watkins, 1998), logistic loss (Ohn Aldrich, 1997), generalized cross-entropy (Zhang & Sabuncu, 2018), and mean absolute error loss (Ghosh et al., 2017). The corresponding true error for $\Phi_{01}^{u}$ is defined as $\mathcal{E}_{\Phi_{01}^{u}}(h)=\mathbb{E}_{(x,y)}[\Phi_{01}^{u}(h,x,y)]$, with its optimal value expressed as $\mathcal{E}^{\ast}_{\Phi_{01}^{u}}(\mathcal{H})=\inf_{h\in\mathcal{H}}\mathcal{E}_{\Phi_{01}^{u}}(h)$.
A key property of a surrogate loss is Bayes-consistency, which ensures that minimizing the surrogate excess risk leads to minimizing the true excess risk (Zhang, 2002; Bartlett et al., 2006; Steinwart, 2007; Tewari & Bartlett, 2007). Formally, $\Phi_{01}^{u}$ is Bayes-consistent with respect to $\ell_{01}$ if, for any sequence $\{h_{k}\}_{k\in\mathbb{N}}\subset\mathcal{H}$, the following implication holds:
@@
This property typically assumes $\mathcal{H}=\mathcal{H}_{\text{all}}$, which may not hold for restricted hypothesis classes such as $\mathcal{H}_{\text{lin}}$ or $\mathcal{H}_{\text{ReLU}}$ (Long & Servedio, 2013; Awasthi et al., 2022; Mao et al., 2024a). To characterize consistency with a particular hypothesis set, Awasthi et al. (2022) introduced $\mathcal{H}$-consistency bounds, which rely on a non-decreasing function $\Gamma:\mathbb{R}^{+}\to\mathbb{R}^{+}$ and take the following form:
@@
Adversarial robustness.
Adversarially robust classification aims to train classifiers that are robust to small, imperceptible perturbations of the input (Goodfellow et al., 2014; Madry et al., 2017). The objective is to minimize the true multiclass loss $\ell_{01}$ evaluated on an adversarial input $x^{\prime}=x+\delta$ (Gowal et al., 2020; Awasthi et al., 2022). A perturbation $\delta$ is constrained by its magnitude, and we define the adversarial region around $x$ as $B_{p}(x,\gamma)=\{x^{\prime}\mid\|x^{\prime}-x\|_{p}\leq\gamma\}$, where $\|\|_{p}$ is the $p$-norm and $\gamma\in(0,1)$ specifies the maximum allowed perturbation.
Definition 3.1 (Adversarial True Multiclass Loss).
The adversarial true multiclass loss ~$\widetilde{\ell}_{01}:\mathcal{H}\times\mathcal{X}\times\mathcal{Y}\to\{0,1\}$ is given by:
~$\widetilde{\ell}_{01}(h,x,y)=\sup_{x^{\prime}\in B_{p}(x,\gamma)}\ell_{01}(h(x^{\prime}),y).$
Similarly to classification, minimizing ~$\widetilde{\ell}_{01}$ is computationally infeasible (Zhang & Agarwal, 2020; Bartlett et al., 2006; Awasthi et al., 2022). To address this, we introduce the following surrogates.
Definition 3.2 (Adversarial Margin Surrogates).
The family of adversarial margin surrogate losses $\widetilde{\Phi}^{\rho,u}_{01}$ from the comp-sum $\rho$-margin family, which approximate the adversarial true multiclass loss ~$\widetilde{\ell}_{01}$ is defined as:
@@
Here, $\Psi^{u}$ and $\Psi_{\rho}$ represent transformations that characterize the behavior of the family. The piecewise-linear and non-convex transformation is defined as $\Psi_{\rho}(v)=\min\left\{\max\left(0,1-\frac{v}{\rho}\right),1\right\}$. Recent studies have shown that algorithms based on smooth, regularized variants of the comp-sum $\rho$-margin losses achieve $\mathcal{H}$-consistency, thereby offering strong theoretical guarantees (Awasthi et al., 2022, 2023; Mao et al., 2023b).
Two-Stage Learning-to-Defer.
The Learning-to-Defer (L2D) framework assigns each input $x\in\mathcal{X}$ to the most cost-effective agent, leveraging the strengths of multiple agents to improve task performance. Agents consist of a fixed primary model and a set of $J$ offline experts. We denote the full agent set as $\mathcal{A}=\{0\}\cup[J]$, where $j=0$ refers to the primary model $g$ defined in Section 3. Each expert $\mathrm{M}_{j}$ outputs a prediction pair $m_{j}(x)=(m_{j}^{h}(x),m_{j}^{f}(x))$, with $m_{j}^{h}(x)\in\mathcal{Y}$ a categorical prediction and $m_{j}^{f}(x)\in\mathcal{T}$ a regression output. The complete set of expert predictions is denoted $m(x)=\big{(}m_{1}(x),\dots,m_{J}(x)\big{)}\in\mathcal{M}$. In the two-stage setting, all agents are trained offline, and only the allocation function is learned, keeping agent parameters fixed.
A rejector $r\in\mathcal{R}$, defined as $r:\mathcal{X}\times\mathcal{A}\to\mathbb{R}$, is trained to select the most suitable agent by computing $r(x)=\arg\max_{j\in\mathcal{A}}r(x,j)$, as proposed by Mao et al. (2024c, 2023a); Montreuil et al. (2024).
Definition 3.3 (Two-Stage L2D Losses).
Let $x\in\mathcal{X}$ and $r\in\mathcal{R}$. The true deferral loss is defined as:
$\ell_{\text{def}}(r,g,m,z)=\sum_{j=0}^{J}c_{j}(g(x),m_{j}(x),z)1_{r(x)=j}.$
Its convex, non-negative, upper-bound surrogate family is given by:
@@
where $c_{j}$ denotes the cost of assigning query $x$ to agent $j\in\mathcal{A}$ (Madras et al., 2018). If $r(x)=0$, the decision is handled by the primary model $g$, which returns prediction $g(x)=(h(w(x)),f(w(x)))$ and incurs cost $c_{0}(g(x),z)=\psi(g(x),z)$, where $\psi(\cdot,\cdot)\in\mathbb{R}^{+}$ is a general loss function.
If $r(x)=j$ for some $j>0$, the query is deferred to expert $j$, incurring a cost $c_{j}(m_{j}(x),z)=\psi(m_{j}(x),z)+\beta_{j}$, where $\beta_{j}\geq 0$ is the consultation cost. The complementary cost function used in the surrogate is:
$\tau_{j}(g(x),m(x),z)=\sum_{i=0}^{J}c_{i}(g(x),m_{i}(x),z)1_{i\neq j},$
@@
which matches the formulation from Mao et al. (2024c); Montreuil et al. (2024). In the case of classification, $\psi$ corresponds to the standard 0–1 loss.
We illustrate how Learning-to-Defer works at inference in Figure 1.
4 Novel Adversarial Attacks on L2D
Motivation.